Any help would be appreciated!

Find the perimeter and area of the rectangle with sides that measure 1 1/2 ft by 4 1/3 ft. State units with your answers.

Answer by checkley79(3341)

P=2L+2W
P=2*1.5+2*4.33
P=3+8.66
P=11.66 FT. FOR THE PERIMETER.
A=LW
A=1.5*4.33
A=6.495 FT^2 ANS. FOR THE AREA.